# Stochastic Optimization with Optimal Importance Sampling

### v2

@@ 2.2 Constrained Stochastic Optimization
We now consider the general constrained stochastic optimization problem. Following [bertsekas2009convex, Proposition 5.4.7], the necessary and sufficient first-order optimality conditions of problem (SO) under Assumption 2.1(i) and 2.1(ii) reduce to:
@@
The previous fact implies that we are looking for a point $\theta^{\star}$ in the affine subspace associated with the active constraints, for which the gradient is orthogonal to the active constraints; see Figure 1(a). This geometric observation motivates all active set methods [nocedal1999numerical], with the simplex method as the primary example.
@@
In the presence of constraints, the performance analysis of stochastic approximation algorithms is however much more delicate. Through a local asymptotic minimax argument, duchi2021asymptotic show that any sequence of approximations $\{\bar{\theta}_{n}\}_{n\in\mathbb{N}}$ which satisfies the CLT
@@
with $\text{Q}:=\text{P}_{A^{\star}_{a}}\nabla^{2}f(\theta^{\star})\text{P}_{A^{\star}_{a}}$, enjoys in fact optimal asymptotic convergence. It can be remarked here that the performance lower bound in equation (8) coincides with the asymptotic variance in (4) of the classical PR-SA applied to the unconstrained problem $\min_{\theta}\,{f}(\theta_{a}+\text{P}_{A^{\star}_{a}}\theta)$, for an arbitrary $\theta_{a}$ satisfying $A^{\star}_{a}\theta_{a}-b^{\star}_{a}=0$. However, the classical PR-SA applied to the unconstrained problem is not a viable algorithm as the active constraints are not known a priori. To solve this issue, duchi2021asymptotic introduce the following variant of Nesterov’s dual averaging (NDA) [nesterov2009primal] iterates
@@
with $\theta_{0}\in\mathbb{R}^{s}$ and $\alpha_{n}=\alpha/n^{\gamma}$, for $\gamma\in(1/2,1)$ and some constant $\alpha>0$. Notice that if the stochastic optimization problem (SO) is unconstrained, then iteration (9) simply reduces to the classical PR-SA. Hence, the iterates $\bar{\theta}_{n}$ in equation (9) can be considered a direct generalization of the classical PR-SA to the constrained setting. Under standard regularity assumptions (similar in nature to those in polyak1992acceleration) and under the condition that the constraint set $\Theta$ is polytopic (as in problem (SO)), duchi2021asymptotic show that the averaged iterates satisfy the almost sure convergence $\bar{\theta}_{n}\overset{\text{a.s.}}{\to}\theta^{\star}$, as well as the CLT (8). Remarkably, the proposed iteration identifies the active constraints in finite time. That is, there exists some (random) finite $N\in\mathbb{N}$ such that $A^{\star}_{a}\theta_{n}-b^{\star}_{a}=0$ and $A^{\star}_{i}\theta_{n}-b_{i}<0$ for all $n\geq N$, and the iterations are identical to those of the PR-SA method in the affine subspace $\{\theta:\;A^{\star}_{a}\theta-b^{\star}_{a}=0\}$ for all $n\geq N$.
@@
(with subsequent averaging for PR-SA) fails to identify the active constraints in finite time. Indeed, there are many instances where the iterates do not satisfy the active constraints with constant non-zero probability at each iteration, and consequently jump off the constraint infinitely often (see [lee2012manifold, duchi2021asymptotic] for details). Nonetheless, davis2023asymptotic show that this standard projected version also achieves the guarantee (8) with minimal asymptotic variance. Moreover, unlike NDA, the optimality of projected RM-SA or PR-SA iterations is not restricted to polytopic constraint sets $\Theta$. As will become clear later, the finite-time identification of the active constraints is a key step in the convergence analysis of our method. For this reason, in this paper we will adopt the NDA iteration (9). $\clubsuit$
@@
The SAA procedure is asymptotically optimal as well, enjoying the CLT (8) (see [shapiro1989asymptotic, Theorem 3.3]). However, for modern large scale problems, online stochastic gradient methods are generally preferred due to their superior computational efficiency. Indeed, equation (12) still requires the solution of an optimization problems whose complexity grows with increasing sample size. $\clubsuit$
@@ 3 Importance Sampling
As discussed earlier, stochastic gradient methods are based on the definition $G(\theta,X):=\nabla_{\theta}F(\theta,X),$ so that $\nabla f(\theta)=\mathbb{E}_{X\sim\mathbb{P}}\left[G(\theta,X)\right]$. However, such stochastic gradients may have very large variance, leading to practical inefficiency. This is often the case in optimization problems involving rare events, where the expected loss to be minimized may be determined by extreme events that occur infrequently, but that are associated with very large costs. With naive Monte-Carlo simulation, such events are infrequently observed, and hence the associated stochastic gradient estimators display high variance, and consequently poor practical performance.
Importance sampling (IS) is a variance reduction technique [glasserman2004monte] which can be employed in the selection of better stochastic gradients. Such technique is based on the following observation:
@@
where ${\mathrm{d}\mathbb{P}}/{\mathrm{d}\mathbb{P}_{\operatorname*{IS}}}:\mathcal{X}\to\mathbb{R}_{+}$ represents the Radon-Nicodym derivative of the original distribution $\mathbb{P}$ with respect to the importance sampling distribution $\mathbb{P}_{\operatorname*{IS}}$. We remark that this derivative exists as long as $\mathbb{P}\ll\mathbb{P}_{\operatorname*{IS}}$ [nikodym1930sur], and unlike Bardou2009computing we do not require the two distributions to be absolutely continuous with respect to the Lebesgue measure (i.e., admit a density function on $\mathbb{R}^{r}$). In the particular case where the support set $\mathcal{X}$ of $\mathbb{P}$ is finite (e.g., in empirical risk minimization problems), the absolute continuity requirement reduces to the condition $\mathbb{P}_{\operatorname*{IS}}(x)>0$ for all $x\in\mathcal{X}$. In view of Lemma 2.4 and Equation (11), an IS distribution $\mathbb{P}_{\operatorname*{IS}}$ is good if
@@
and stochastic gradients by $G_{\mu}(\theta,x):=\ell(x,\mu)G(\theta,x)$. Moreover, for the IS class to be well behaved, we impose the following assumptions on the function $\ell$, which are similar to those considered in [Bardou2009computing].
@@ 3.1 Exponential Tilting
In certain cases, the IS distribution $\mathbb{P}_{\mu}$ belongs to the same parametric family as $\mathbb{P}$. This is the case for example when the original density belongs to the exponential family of distributions (e.g., normal, exponential, Poisson, chi-squared, etc.), simplifying random variable generation during Monte Carlo simulations. The exponential tilting has also proven to be fundamental in the context of rare-event simulation and large deviation theory, where often it is the unique efficient simulation distribution choice (see [bucklew2004introduction, Chapter $5.2$]). The following lemma states that Assumption 3.1(ii) is also satisfied. Its proof follows immediately from Lebesgue’s dominated convergence, and is therefore omitted.
@@ 3.2 Mean Translation
can be evaluated efficiently. Such translations of the density function are generally used to place more probability mass in a rare event region, and have been successfully employed in the context of digital communication systems [bucklew2004introduction].
@@ 3.3 Mixture Models
Finally, consider a setting in which the decision-maker has access to $I\in\mathbb{N}$ distinct importance samplers. That is, we have $G_{i}(\theta,X)=G(\theta,X)\ell_{i}(X)$ so that, for all $i\in[I]$, we have $\nabla f(\theta)=\mathbb{E}_{\mathbb{P}_{i}}\left[G_{i}(\theta,X)\right]$ and where we assume that we can sample from each distribution $\mathbb{P}_{i}$ efficiently. Given any $\mu\in\mathbb{R}^{I}_{+}$ such that $\sum_{i=1}^{I}\mu_{i}=1$, we can consider the mixture distribution $\mathbb{P}_{\mu}=\sum_{i=1}^{I}\mu_{i}\mathbb{P}_{i}$ from which we can sample efficiently given access to samples from each of the IS distributions $\mathbb{P}_{i}$, for $i\in[I]$. Such mixture models are considered for instance in [bucklew2004introduction, Section 5.2.2] and are successfully employed to define efficient important samplers in the context of a large deviation principle.
@@ 4 Adaptive Importance Sampling
Given only black-box access to the optimization problem (SO), the optimizer has access only to samples from $G(\theta,X)$, where $G(\theta,X):=\nabla_{\theta}F(\theta,X)$ and $\nabla f(\theta)=\mathbb{E}_{X\sim\mathbb{P}}\!\left[G(\theta,X)\right]$. Under this information structure, the NDA iteration proposed in [duchi2021asymptotic] is optimal in the sense discussed in Section 2. In this paper we go beyond the black-box model and assume that sampling from an IS class is a viable option.
@@
Using the notation introduced in equation (13) and below in equation (18), the two conditions in Assumption 4.1(iii) are equivalent to $\mathbb{E}_{X^{(\mu)}\sim\mathbb{P}_{\mu}}\left[\,\|G_{\mu}(\theta,X^{(\mu)})\|^{2}\right]\leq G^{2}_{M}<\infty$ and $\mathbb{E}_{X\sim\mathbb{P}}\left[\,\|H(\theta,\mu,X)\|^{2}\right]\leq H^{2}_{M}<\infty$, respectively. Such a second-moment boundedness assumption on the gradients is standard in the analysis of stochastic approximation algorithms; see, e.g., [nemirovski2009robust]. In our setting, this condition is often satisfied due to the compactness of the sets $\Theta$ and $\mathcal{M}$, for many common choices of the distribution $\mathbb{P}$. As an example, it can be easily checked that this is the case if $\mathbb{P}$ is Gaussian and the parameter $\mu$ comes from the exponential tilting IS class.
@@ 4.1 Secondary Importance Sampling
    First, in particular settings, a natural choice of secondary IS parameters can be available. In the context of exponential tilting discussed in Section 3.1, with $\mathcal{M}=$-$\mathcal{M}$, lemaire2010unconstrained suggest using the same exponential tilting class for the secondary IS class together with the simple choice $\nu_{k}=-\mu_{k}$, so that
@@
    This choice has been exploited in [lemaire2010unconstrained] to bound and control the growth of the stochastic gradient in order to ensure the almost sure convergence of the adaptive IS procedure even when the set $\mathcal{M}$ is not bounded.
@@
    Secondly, as suggested by [he2023adaptive], we may assume that there exists a readily available good importance sampler in the secondary importance sampling class $\left\{\mathbb{P}_{\nu}:\nu\in\mathcal{V}\right\}$ for estimating the gradient $\nabla_{\mu}\mathbb{E}_{X\sim\mathbb{P}}[V(\theta,\mu,X)]$, for a given $\theta$ and $\mu$. That is, we may assume to have access to a mapping $I$ so that $\nu=I(\theta,\mu)$ is associated with an asymptotic variance $\text{Var}_{X^{(\nu)}\sim\mathbb{P}_{\nu}}\left[H_{\nu}(\theta,\mu,X^{(\nu)})\right]$ that is sufficiently small. In certain cases, large deviations theory can be used to derive such a mapping.
@@ 5.1 Almost Sure Convergence
    There exists $c_{2}<\infty$ such that for all $\theta\in\Theta$ and $\mu\in\mathcal{M}$:
@@
Term 1: We start with the maximization over $\theta$. Defining $\ell_{1}(\theta):=\frac{1}{2}\|\theta\|^{2}+\delta_{\Theta}(\theta)$, we have that the maximum over $\theta$ is precisely the convex conjugate $\ell_{1}^{*}(-\sum_{k=0}^{n}\alpha_{k+1}G_{k})$. As $\min_{\theta\in\mathbb{R}^{s}}\{\langle\sum_{k=0}^{n}\alpha_{k+1}G_{k},\theta\rangle+\frac{1}{2}\|\theta\|^{2}+\delta_{\Theta}(\theta)\}$ is the Moreau envelope of $\sum_{k=0}^{n}\alpha_{k+1}G_{k}+\delta_{\Theta}(\theta)$ evaluated at zero, we have that the gradient $\nabla\ell_{1}^{*}(-\sum_{k=0}^{n}\alpha_{k+1}G_{k})$ is equal to proximal map evaluated at zero [rockafellar2009variational, Theorem 2.26], i.e.,
@@
Proving the almost sure converge of iterates in SA algorithms is by now a relatively standard procedure, which generally employs the use of results such as Theorem 1 in robbins1971convergence (see Lemma A.1). Interestingly, the joint NDA iteration (19) introduces a technical challenge that is generally not present in the existing proofs in the literature. We detail this in the following remark.
@@
For the joint NDA iteration (19), the standard choice in the literature for $R_{n}$ (see, e.g., [duchi2021asymptotic, Theorem 2]) leads to the following term as a natural candidate for $C_{n}$:

### v3

@@ 2.2 Constrained Stochastic Optimization
We now consider the general constrained stochastic optimization problem. Following [Bertsekas, 2009, Proposition 5.4.7], the necessary and sufficient first-order optimality conditions of problem (SO) under Assumption 2.1(i) and 2.1(ii) reduce to:
@@
The previous fact implies that we are looking for a point $\theta^{\star}$ in the affine subspace associated with the active constraints, for which the gradient is orthogonal to the active constraints; see Figure 1(a). This geometric observation motivates all active set methods Nocedal and Wright [1999], with the simplex method as the primary example.
@@
In the presence of constraints, the performance analysis of stochastic approximation algorithms is however much more delicate. Through a local asymptotic minimax argument, Duchi and Ruan [2021, Theorems 1 and 4] show that any sequence of approximations $\{\bar{\theta}_{n}\}_{n\in\mathbb{N}}$ which satisfies the CLT
@@
with $\text{Q}:=\text{P}_{A^{\star}_{a}}\nabla^{2}f(\theta^{\star})\text{P}_{A^{\star}_{a}}$, enjoys in fact optimal asymptotic convergence. It can be remarked here that the performance lower bound in equation (8) coincides with the asymptotic variance in (4) of the classical PR-SA applied to the unconstrained problem $\min_{\theta}\,{f}(\theta_{a}+\text{P}_{A^{\star}_{a}}\theta)$, for an arbitrary $\theta_{a}$ satisfying $A^{\star}_{a}\theta_{a}-b^{\star}_{a}=0$. However, the classical PR-SA applied to the unconstrained problem is not a viable algorithm as the active constraints are not known a priori. To solve this issue, Duchi and Ruan [2021] introduce the following variant of Nesterov’s dual averaging (NDA) Nesterov [2009] iterates
@@
with $\theta_{0}\in\mathbb{R}^{s}$ and $\alpha_{n}=\alpha/n^{\gamma}$, for $\gamma\in(1/2,1)$ and some constant $\alpha>0$. Notice that if the stochastic optimization problem (SO) is unconstrained, then iteration (9) simply reduces to the classical PR-SA. Hence, the iterates $\bar{\theta}_{n}$ in equation (9) can be considered a direct generalization of the classical PR-SA to the constrained setting. Under standard regularity assumptions (similar in nature to those in Polyak and Juditsky [1992]) and under the condition that the constraint set $\Theta$ is polytopic (as in problem (SO)), Duchi and Ruan [2021] show that the averaged iterates satisfy the almost sure convergence $\bar{\theta}_{n}\overset{\text{a.s.}}{\to}\theta^{\star}$, as well as the CLT (8). Remarkably, the proposed iteration identifies the active constraints in finite time. That is, there exists some (random) finite $N\in\mathbb{N}$ such that $A^{\star}_{a}\theta_{n}-b^{\star}_{a}=0$ and $A^{\star}_{i}\theta_{n}-b_{i}<0$ for all $n\geq N$, and the iterations are identical to those of the PR-SA method in the affine subspace $\{\theta:\;A^{\star}_{a}\theta-b^{\star}_{a}=0\}$ for all $n\geq N$.
@@
(with subsequent averaging for PR-SA) fails to identify the active constraints in finite time. Indeed, there are many instances where the iterates do not satisfy the active constraints with constant non-zero probability at each iteration, and consequently jump off the constraint infinitely often (see Lee et al. [2012]; Duchi and Ruan [2021] for details). Nonetheless, Davis et al. [2023] show that this standard projected version also achieves the guarantee (8) with minimal asymptotic variance. Moreover, unlike NDA, the optimality of projected RM-SA or PR-SA iterations is not restricted to polytopic constraint sets $\Theta$. As will become clear later, the finite-time identification of the active constraints is a key step in the convergence analysis of our method. For this reason, in this paper we will adopt the NDA iteration (9). $\clubsuit$
@@
The SAA procedure is asymptotically optimal as well, enjoying the CLT (8) (see [Shapiro, 1989, Theorem 3.3]). However, for modern large scale problems, online stochastic gradient methods are generally preferred due to their superior computational efficiency. Indeed, equation (12) still requires the solution of an optimization problems whose complexity grows with increasing sample size. $\clubsuit$
@@ 3 Importance Sampling
As discussed earlier, stochastic gradient methods are based on the definition $G(\theta,X):=\nabla_{\theta}F(\theta,X),$ so that $\nabla f(\theta)=\mathbb{E}_{X\sim\mathbb{P}}\left[G(\theta,X)\right]$. However, such stochastic gradients may have very large variance, leading to practical inefficiency. This is often the case in optimization problems involving rare events, where the expected loss to be minimized may be determined by extreme events that occur infrequently, but that are associated with very large costs. With naïve Monte-Carlo simulation, such events are infrequently observed, and hence the associated stochastic gradient estimators display high variance, and consequently poor practical performance.
Importance sampling (IS) is a variance reduction technique [Glasserman, 2004] which can be employed in the selection of better stochastic gradients. Such technique is based on the following observation:
@@
where ${\mathrm{d}\mathbb{P}}/{\mathrm{d}\mathbb{P}_{\operatorname*{IS}}}:\mathcal{X}\to\mathbb{R}_{+}$ represents the Radon-Nicodym derivative of the original distribution $\mathbb{P}$ with respect to the importance sampling distribution $\mathbb{P}_{\operatorname*{IS}}$. We remark that this derivative exists as long as $\mathbb{P}\ll\mathbb{P}_{\operatorname*{IS}}$ Nikodym [1930], and unlike Bardou et al. [2009] we do not require the two distributions to be absolutely continuous with respect to the Lebesgue measure (i.e., admit a density function on $\mathbb{R}^{r}$). In the particular case where the support set $\mathcal{X}$ of $\mathbb{P}$ is finite (e.g., in empirical risk minimization problems), the absolute continuity requirement reduces to the condition $\mathbb{P}_{\operatorname*{IS}}(x)>0$ for all $x\in\mathcal{X}$. In view of Lemma 2.4 and Equation (11), an IS distribution $\mathbb{P}_{\operatorname*{IS}}$ is good if
@@
and stochastic gradients by $G_{\mu}(\theta,x):=\ell(x,\mu)G(\theta,x)$. Moreover, for the IS class to be well behaved, we impose the following assumptions on the function $\ell$, which are similar to those considered in [Bardou et al., 2009].
@@ 3.1 Exponential Tilting
In certain cases, the IS distribution $\mathbb{P}_{\mu}$ belongs to the same parametric family as $\mathbb{P}$. This is the case for example when the original density belongs to the exponential family of distributions (e.g., normal, exponential, Poisson, chi-squared, etc.), simplifying random variable generation during Monte Carlo simulations. The exponential tilting has also proven to be fundamental in the context of rare-event simulation and large deviation theory, where often it is the unique efficient simulation distribution choice (see [Bucklew and Bucklew, 2004, Chapter $5.2$]). The following lemma states that Assumption 3.1(ii) is also satisfied. Its proof follows immediately from Lebesgue’s dominated convergence, and is therefore omitted.
@@ 3.2 Mean Translation
can be evaluated efficiently. Such translations of the density function are generally used to place more probability mass in a rare event region, and have been successfully employed in the context of digital communication systems Bucklew and Bucklew [2004].
@@ 3.3 Mixture Models
Finally, consider a setting in which the decision-maker has access to $I\in\mathbb{N}$ distinct importance samplers. That is, we have $G_{i}(\theta,X)=G(\theta,X)\ell_{i}(X)$ so that, for all $i\in[I]$, we have $\nabla f(\theta)=\mathbb{E}_{\mathbb{P}_{i}}\left[G_{i}(\theta,X)\right]$ and where we assume that we can sample from each distribution $\mathbb{P}_{i}$ efficiently. Given any $\mu\in\mathbb{R}^{I}_{+}$ such that $\sum_{i=1}^{I}\mu_{i}=1$, we can consider the mixture distribution $\mathbb{P}_{\mu}=\sum_{i=1}^{I}\mu_{i}\mathbb{P}_{i}$ from which we can sample efficiently given access to samples from each of the IS distributions $\mathbb{P}_{i}$, for $i\in[I]$. Such mixture models are considered for instance in [Bucklew and Bucklew, 2004, Section 5.2.2] and are successfully employed to define efficient important samplers in the context of a large deviation principle.
@@ 4 Adaptive Importance Sampling
Given only black-box access to the optimization problem (SO), the optimizer has access only to samples from $G(\theta,X)$, where $G(\theta,X):=\nabla_{\theta}F(\theta,X)$ and $\nabla f(\theta)=\mathbb{E}_{X\sim\mathbb{P}}\!\left[G(\theta,X)\right]$. Under this information structure, the NDA iteration proposed in Duchi and Ruan [2021] is optimal in the sense discussed in Section 2. In this paper we go beyond the black-box model and assume that sampling from an IS class is a viable option.
@@
Using the notation introduced in equation (13) and below in equation (18), the two conditions in Assumption 4.1(iii) are equivalent to $\mathbb{E}_{X^{(\mu)}\sim\mathbb{P}_{\mu}}\left[\,\|G_{\mu}(\theta,X^{(\mu)})\|^{2}\right]\leq G^{2}_{M}<\infty$ and $\mathbb{E}_{X\sim\mathbb{P}}\left[\,\|H(\theta,\mu,X)\|^{2}\right]\leq H^{2}_{M}<\infty$, respectively. Such a second-moment boundedness assumption on the gradients is standard in the analysis of stochastic approximation algorithms; see, e.g., Nemirovski et al. [2009]. In our setting, this condition is often satisfied due to the compactness of the sets $\Theta$ and $\mathcal{M}$, for many common choices of the distribution $\mathbb{P}$. As an example, it can be easily checked that this is the case if $\mathbb{P}$ is Gaussian and the parameter $\mu$ comes from the exponential tilting IS class.
@@ 4.1 Secondary Importance Sampling
    First, in particular settings, a natural choice of secondary IS parameters can be available. In the context of exponential tilting discussed in Section 3.1, with $\mathcal{M}=$-$\mathcal{M}$, Lemaire and Pagès [2010] suggest using the same exponential tilting class for the secondary IS class together with the simple choice $\nu_{k}=-\mu_{k}$, so that
@@
    This choice has been exploited in Lemaire and Pagès [2010] to bound and control the growth of the stochastic gradient in order to ensure the almost sure convergence of the adaptive IS procedure even when the set $\mathcal{M}$ is not bounded.
@@
    Secondly, as suggested by He et al. [2023], we may assume that there exists a readily available good importance sampler in the secondary importance sampling class $\left\{\mathbb{P}_{\nu}:\nu\in\mathcal{V}\right\}$ for estimating the gradient $\nabla_{\mu}\mathbb{E}_{X\sim\mathbb{P}}[V(\theta,\mu,X)]$, for a given $\theta$ and $\mu$. That is, we may assume to have access to a mapping $I$ so that $\nu=I(\theta,\mu)$ is associated with an asymptotic variance $\text{Var}_{X^{(\nu)}\sim\mathbb{P}_{\nu}}\left[H_{\nu}(\theta,\mu,X^{(\nu)})\right]$ that is sufficiently small. In certain cases, large deviations theory can be used to derive such a mapping.
@@ 5.1 Almost Sure Convergence
    There exists $c_{2},c_{3}<\infty$ such that for all $\theta\in\Theta$ and $\mu\in\mathcal{M}$:
@@
Term 1: We start with the maximization over $\theta$. Defining $\ell_{1}(\theta):=\frac{1}{2}\|\theta\|^{2}+\delta_{\Theta}(\theta)$, we have that the maximum over $\theta$ is precisely the convex conjugate $\ell_{1}^{*}(-\sum_{k=0}^{n}\alpha_{k+1}G_{k})$. As $\min_{\theta\in\mathbb{R}^{s}}\{\langle\sum_{k=0}^{n}\alpha_{k+1}G_{k},\theta\rangle+\frac{1}{2}\|\theta\|^{2}+\delta_{\Theta}(\theta)\}$ is the Moreau envelope of $\sum_{k=0}^{n}\alpha_{k+1}G_{k}+\delta_{\Theta}(\theta)$ evaluated at zero, we have that the gradient $\nabla\ell_{1}^{*}(-\sum_{k=0}^{n}\alpha_{k+1}G_{k})$ is equal to proximal map evaluated at zero [Rockafellar and Wets, 2009, Theorem 2.26], i.e.,
@@
Proving the almost sure converge of iterates in SA algorithms is by now a relatively standard procedure, which generally employs the use of results such as Theorem 1 in Robbins and Siegmund [1971] (see Lemma A.1). Interestingly, the joint NDA iteration (19) introduces a technical challenge that is generally not present in the existing proofs in the literature. We detail this in the following remark.
@@
For the joint NDA iteration (19), the standard choice in the literature for $R_{n}$ (see, e.g., [Duchi and Ruan, 2021, Theorem 2]) leads to the following term as a natural candidate for $C_{n}$: